Outline

Contents

1  1
2  1
2.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1

Make Titles Informative. Make Titles Informative. Make Titles Informative.

2

Make Titles Informative. Make Titles Informative. Make Titles Informative.

2.1

Make Titles Informative. Make Titles Informative. Make Titles Informative. License covering this document

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005 Nick Urbanik <nicku@nicku.org>

You can redistribute modified or unmodified copies of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation — either version 2 of the License or (at your option) any later version.